Please help me I having trouble with this problem.
Complete the ordered pairs so that each is a solution for the given equations.
3x+4y=12 (0, ),(3/4),( ,0),(8/3)

Answer by malakumar_kos@yahoo.com(315)


Complete the ordered pairs so that each is a solution for the given equations.
3x+4y=12 (0, ),(3/4),( ,0),(8/3)
equation is 3x+4y = 12
1) (0,) When x =0 then 3(0)+4y = 12 or 4y = 12 or y = 12/4 = 3
2)(3/4,) When x = 3/4 then 3(3/4)+4y = 12 4y = 12-9/4 = 48-9/4 = 39/4
y = 39/16
3)( ,0) When y = 0 then 3x+4(0) = 12 or 3x = 12 x = 12/3 = 4

4)(8/3, ) When x = 8/3 then 3(8/3)+4y = 12 or 4y = 12-8 = 4 or y = 1
The required ordered pairs are (0,3),(3/4,39/16),(4,0),(8/3,1)